### v1

# AI-boosted rare event sampling to characterize extreme weather

Abstract
Assessing the frequency and intensity of extreme weather events, and understanding how climate change affects them, is crucial for developing effective adaptation and mitigation strategies [1, 2, 3, 4, 5]. However, observational datasets are too short and physics-based global climate models (GCMs) are too computationally expensive to obtain robust statistics for the rarest, yet most impactful, extreme events [6]. AI-based emulators have shown promise for predictions at weather and even climate timescales [7, 8, 9, 10, 11, 12, 13], but they struggle on extreme events with few, or no, examples in their training dataset [14, 15, 16, 17]. Rare event sampling (RES) algorithms have previously demonstrated success for some extreme events [18, 19, 20, 21, 22, 23], but their performance depends critically on a hard-to-identify “score function”, which guides efficient sampling by a GCM. Here, we develop a novel algorithm, AI+RES, which uses ensemble forecasts of an AI weather emulator [7, 8, 9, 10] as the score function to guide highly efficient resampling of the GCM and generate robust (physics-based) extreme weather statistics and associated dynamics at $30-300\times$ lower cost. We demonstrate AI+RES on mid-latitude heatwaves, a challenging test case requiring a score function with predictive skill many days in advance. AI+RES, which synergistically integrates AI, RES, and GCMs, offers a powerful, scalable tool for studying extreme events in climate science, as well as other disciplines in science and engineering where rare events and AI emulators are active areas of research.
Introduction
Rare, extreme weather and climate events have enormous economic and social costs [24, 25], but it is difficult to estimate their frequency and potential impact from observational data or high-fidelity physics-based GCM simulations. Extreme value theory [26] can be used to extrapolate from short historical records, but it is often overwhelmed by uncertainty [27, 28, 6, 22], and does not provide examples of the event of interest for physical interrogation. GCMs require simulation lengths at least ten times longer than the return time of the event of interest for accurate sampling, which is not feasible given the computational cost of state-of-the-art physics-based GCMs to fully resolve all relevant scales and processes. In response to this need, rare event sampling (RES) and artificial intelligence (AI) emulator techniques have been developed, but both have proven insufficient on their own to fully resolve the problem.
RES is a set of importance sampling tools that focus computation on the rare event of interest [18, 19, 29, 20, 30, 22, 23]. RES entails scheduled duplication (“splitting”) or termination (“killing”) of members of an ensemble of model simulations, that otherwise evolve independently, to promote efficient sampling of the targeted rare event (Fig. 1a). The probability that any one ensemble member (a trajectory) is duplicated or terminated is determined by a user-chosen score function that can be evaluated for any trajectory at the scheduled resampling times. Despite promising results, the full power of RES has not been realized for climate and weather due primarily to the difficulty in identifying effective score functions. Choosing a score function typically requires extensive domain knowledge and a process of trial and error with associated costs in computation and time that may swamp any advantages of using RES. Standard RES uses simple persistence as a score function: it relies on the current value of the index variable (which defines the extreme event of interest; e.g., 5-day mean temperature), rather than its value at the time of the event. This choice of score function has proven highly effective for sampling extremes of long-term averages, e.g., seasonal means [18, 19, 21, 22, 23].
However, this approach fails catastrophically when the current value of the index variable is a poor predictor of its future value [19, 20], as is the case for many phenomena in complex atmospheric dynamics; examples include blocking-driven heatwaves, which have led to some of the most socio-economically impactful extreme events in recent memory in France, Russia, and U.S. Pacific Northwest [31, 32, 33]. Previous work has therefore established the strong potential of RES, but only if an efficient method for finding an effective score function can be identified. Recently, a set of interesting papers studied unprecedented extreme events using ensemble boosting, a cousin of RES [34, 35]. In its initial formulation [34], ensemble boosting could not be used to estimate event probabilities. In a more recent version [35, 36], however, probabilities were estimated by adopting a rare event simulation approach with a one-step splitting scheme. This similar method could benefit from an analogous AI score function.
One of the most exciting recent developments in climate science and scientific AI has been the introduction of auto-regressive AI weather emulators such as FourCastNet, Pangu, GraphCast, and GenCast [7, 8, 9, 10]. AI weather emulators are neural networks that are trained on spatio-temporal data from historical observations or physics-based GCM simulations to advance the state of the system forward in time for a short period (e.g., 6 hours) and then applied recursively to produce forecasts over longer periods. Various analyses have confirmed that AI weather emulators trained on high-resolution observation-derived reanalysis datasets can outperform the best physics-based numerical weather models for short- and medium-range ($\sim$10-15 day) forecasts [9, 8, 37]. More importantly, these emulators are up to $10^{4}$ times faster than state-of-the-art physics-based models [7], suggesting that they might be useful in reducing the uncertainty associated with rare extreme weather events. Specifically, extremely long and/or large ensemble emulations can be generated so that accurate “direct sampling” is possible, even for the rarest events [38, 39]. The problem with this method, however, is that AI emulators are trained on available historical records or expensive high-fidelity GCM simulations that may contain few, or no, examples of the rarest extreme events. Characterizing rare events using direct simulation with an AI emulator would therefore require not only reliable out-of-distribution extrapolation to extreme events beyond the training dataset but also capturing their frequency correctly, which recent work has suggested the AI emulators struggle with [14, 15, 17].
In this paper, we leverage advances in AI weather emulation to build an effective score function for RES, combining recent developments in both areas to construct a robust algorithm to characterize rare, extreme events (Fig. 1). As a demonstration of this new framework (AI+RES), we focus on simulating extreme heat waves and apply our approach to a GCM of intermediate complexity, PlaSim [40], which allows for extensive long-term control runs to compute baseline rare event statistics and rigorously evaluate the performance of our approach. Heat waves are a well-motivated application area because they are the deadliest extreme weather events and are expected to worsen under climate change [41, 42, 43]. We show that AI+RES provides accurate rare event statistics at a numerical speed-up of up to a factor of $\mathcal{O}(100)$, whereas standard RES fails to even yield examples of the rarest events of interest. This paper serves as a demonstration of a novel methodology that can overcome the primary difficulty of RES and unleash its power widely across any field for which an AI emulator can be constructed.
Figure 1: Schematic of the AI+RES framework. We train an AI weather emulator on 100 years of PlaSim GCM data, then use it to guide a RES algorithm. (a) In our proposed method (AI+RES), we run a PlaSim ensemble simulation with NN parallel simulations, called walkers and denoted $X_{t}^{i}$. (b) At each resampling time $t_{k}$ in the algorithm (vertical dashed lines), we perform an ensemble forecast with the AI emulator for each PlaSim walker until the target time $t_{f}$, then duplicate the more promising ones based on these forecasts (Algorithm 1 and Eq. (6)). This allows us to both generate a large catalog of extreme events from physics-based simulations (c) and to compute unbiased probabilities for these rare events (Eq. (4)) (d) at a fraction of the cost of Direct Numerical Simulation (DNS). An example of schematic (a) with actual data from an AI+RES experiment can be found in Extended Data Fig. 3.
AI+RES
Figure 1 outlines the AI+RES algorithm. We use an AI emulator, trained to predict the state of a physical system at time $t+\Delta t$ conditioned on the state at time $t$, to address the major shortcoming in RES algorithms, namely, the choice of a score function $\theta$.
The RES splitting algorithm proceeds as follows: we perform NN parallel simulations (walkers) using a physics-based model (e.g., a GCM), starting at time $t_{0}$ with small initial condition perturbations. At user-specified resampling times, $\{t_{k}\}_{k=1}^{K}$, we selectively duplicate or terminate the $i^{th}$ walker according to the value of the score function $\theta^{i}_{k}$ until the walkers reach the final time horizon $t_{f}=t_{K}$. Walkers with a large score are more likely to be duplicated, while walkers with a small score are more likely to be terminated. After resampling, we continue the GCM simulation of the selected walkers until the next resampling time, when the process is repeated. Each walker is accompanied by a statistical weight $w^{i}_{k}$ that allows unbiased estimation of rare event statistics of interest.
RES applications have most often used simple persistence as the score function: the current value of the index variable (which defines the extreme event of interest, e.g., temperature) is used as a proxy for its future value. In this paper, we introduce a new approach for constructing the score function based on an AI weather emulator. At each resampling time $t_{k}$, and from every current walker state, we perform an ensemble forecast of the system state (with $M$ members) until the final time $t_{f}$ using the AI emulator (see the AI Emulator section in Methods). The score function for each RES walker is then the mean value of the index variable at time $t_{f}$ over the AI emulator ensemble forecast for that walker. To the extent that the AI emulator faithfully reproduces the GCM’s dynamics, the prediction furnished by the AI emulator ensemble is ideally suited for use in constructing the score function. Our results show that the AI emulator’s weather forecast skill is sufficient to provide a very effective score function for AI+RES. More details on our implementation are given in the Methods section, including our choice of hyperparameters for the RES Algorithm 1 in Extended Data Table 2.
Mid-latitude heatwaves
In this paper, we focus on the sampling of rare mid-latitude heatwaves due to their societal impact and the challenge they pose to traditional (standard) RES. Specifically, following previous studies [18, 44], we focus on events characterized by large values of the spatio-temporal average of 2-meter temperature T2m over a geographic region $\mathcal{R}$ during a time interval $[t,t+L]$. This evolving spatio-temporal average is our index variable $A_{L}(t)$:
$A_{L}(t):=\frac{1}{L}\int_{t}^{t+L}\left(\frac{1}{\mathcal{R}}\int_{\mathcal{R}}\text{T}_{2m}(\vec{r},s),d\vec{r}\right)ds.$ (1)
We want to estimate the return times of the events $A_{L}(t_{f})>a$ for large thresholds $a$. We use $L=7$ days, a window size small enough to capture the peak of mid-latitude heatwaves yet large enough to have significant societal impacts. In this study, we consider two regions, $\mathcal{R}$, each defined as a $3\times 3$ grid point domain centered over France and Chicago (Extended Data Fig. 1). We initialize simulations at $t_{0}$=July 2 and integrate forward the PlaSim GCM until $t_{f}+L$, with $t_{f}=$ August 1, chosen to coincide with the climatological peak of $A_{L}(t_{f})$.
Ground truth and baselines
To evaluate the performance of AI+RES, we compare it to a ground truth and five baselines. The ground truth, direct numerical simulation (DNS) with $N=50,000$, is a very large ensemble that is only possible due to the computational efficiency of PlaSim. The first baseline, DNS, $N=400$, is a DNS with PlaSim using the same ensemble size as in the AI+RES algorithm. The second baseline, Standard-RES, uses the RES algorithm but with the traditional persistence score function used in state-of-the-art weather and climate applications [18, 19, 21, 22, 23]. The third baseline, AI-DNS, is a DNS ensemble using the AI emulator started from the same initial conditions as DNS and RES, with the same ensemble size as the ground truth ($N=50,000$). This allows us to determine whether the AI emulator has learned enough from its 100-year training period to extrapolate to tail events in PlaSim. The fourth baseline, EVT, applies Extreme Value Theory, specifically the Generalized Pareto Distribution (GPD) within a Peak-over-Threshold (POT) framework, to PlaSim datasets of the same size as the AI+RES experiments ($N=400$). We estimate uncertainty in this method by performing the fits on 100 different datasets of equal size. The fifth baseline, PFS+RES, yields an upper bound on algorithm performance using a perfect-forecast system as the score function. In this baseline, we use an ensemble forecast generated with PlaSim itself as the score function (Extended Data). However, this baseline is too expensive for more complex GCMs; it is included to show how the algorithm would perform with a “perfect emulator”.
AI+RES accurately estimates long return period events
Figure 2: Return-period curve for the T2m 7-day average over France (left panel) and Chicago (right panel). Black dots are obtained with a $N=50,000$ member DNS and serve as ground truth. The red dots are from DNS, but with the same computational budget as AI+RES ($N=400$). The cyan dots are obtained by running a $N=50,000$ member ensemble with the AI emulator only (AI-DNS). The solid blue line shows the median return-period curve produced by 10 independent realizations of the AI+RES algorithm with $N=400$ walkers and the blue shaded area shows the 10th and 90th percentiles. The solid yellow line shows the median return-period curve produced by 10 independent realizations of the standard RES algorithm with $N=400$ walkers, and the yellow shaded area shows the 10th and 90th percentiles. The gray shaded area (EVT) is obtained by fitting different GPD distributions with 100 independent training datasets of size $N=400$, and showing the 10th and 90th percentiles as confidence interval.
In both France and Chicago, AI+RES produces accurate, unbiased return period estimates up to 50,000 years, despite using ensembles with only $N=400$ walkers; see Fig. 2 (the 50,000-year limit is imposed by the size of our reference ensemble for validating the procedure). In stark contrast, DNS of the same computational cost ($N=400$) can only produce accurate return period estimates up to about 50 years, 1,000 times shorter. This dramatic difference underscores the utility of RES when an effective score function can be identified. Moreover, Standard-RES saturates for return periods longer than about 100 years, falsely suggesting an upper limit on heatwave intensity. This indicates that persistence misses important trajectories to rare heat waves that the AI-based score function is able to identify effectively. The AI-DNS baseline is biased both in the mean and variability of the observable, producing return period estimates that fail even on a decadal timescale (the correct AI-DNS estimate at a return period of about 2,000 years in France is an accident of the curves crossing). It is notable that despite this failure of AI-DNS, the AI emulator’s weather forecast skill can still serve as a useful score function by effectively ranking progress of walkers toward strong heat waves. Finally, the correct return periods do fall within the uncertainty of the EVT baseline even at the longest timescales, but the uncertainty estimate is up to four times larger than that from AI+RES for return periods on the order of 10,000 years. The EVT estimates depend so strongly on the particular years fed into them that they cannot be relied on to produce accurate return period estimates. Also, EVT, unlike AI+RES, does not provide information about the identified extremes and their dynamics.
AI+RES provides a $100\times$ computational speed-up for rare events
Figure 3: Speed-up factors for the AI+RES algorithm over France and Chicago. Left panel: Variance speed-up factor as a function of return period (Eq. (13)). This measures the reduction in the variance of return period estimates compared to DNS (see Methods). Right panel: Sampling speed-up factor as a function of return period (Eq. (10)), averaged over 10 independent algorithm realizations. This measures the gain in the number of extreme samples produced by the algorithm compared to DNS (see Methods).
Now that we have established that AI+RES is unbiased, we can quantify its computational speed-up relative to DNS using efficiency at sampling extreme events (Eq. (10)) and variance of rare probability estimators (Eq. (13)). Both methods tend to yield greater computational gains for rarer events (Fig. 3), with speed-ups of more than 100 for a return period of 10,000 years. The only exception is that the variance speed-up estimate for the France region peaks at a factor of 25 around the 2,500-year return period, then gradually decreases to 10 for the rarest events. This non-monotonic trend is most likely due to a sampling error (see Variance speed-up factor in the Methods section). Moreover, we hypothesize that the difference in variance speed-up for the two regions is due to regional differences in AI emulator weather forecast skill, potentially due to differences in soil moisture variability, as pointed out in Fig. S6 in the SI.
The rare events generated by AI+RES are physically realistic
Figure 4: Composite maps of heatwaves over France with return periods exceeding 100 years. Events are defined by Eq. (1) with $L=$ 3 days. The first row shows daily mean Z500 anomaly composites three days before the heatwave onset; the second row shows the $L$-day average Z500 anomaly composites during the heatwave; the third row shows the $L$-day average T2m anomaly composites during the heatwave. Left column: DNS with N=50, [^0] N=50{,}000 (ground truth). Middle column: results from the AI+RES algorithm with $N=400$. Right column: DNS with $N=400$. See Fig. S5 in the SIfor the same analysis but over Chicago.
In addition to greatly reduced variance in return period estimations, a major advantage of AI+RES over the EVT baseline is that it generates full physical trajectories leading to the rare event of interest. This is particularly valuable for investigating the dynamics of the rare event and its precursors [23].
In Fig. 4, we illustrate this capability by presenting composite maps associated with heatwaves over France with a return period longer than 100 years from the ground-truth DNS (N=50, [^0] N=50{,}000), samples generated with AI+RES ($N=400$), and samples generated from the DNS with the same ensemble ($N=400$). A strong precursor pattern emerges in the ground-truth DNS three days before the heatwave onset, marked by a pronounced positive 500-hPa geopotential height anomaly over the British Isles, flanked by a negative anomaly over the Labrador Sea. A distinct negative anomaly off the coast of Portugal suggests potential influence from a cut-off low, while a weaker anomaly is visible over Eastern Europe. At the heatwave peak, this synoptic pattern shifts: the high-pressure anomaly migrates eastward, the Labrador Sea anomaly propagates into the North Atlantic, and the Eastern European anomaly intensifies. The AI+RES composite shows similar patterns and captures the correct evolution, although it is slightly noisier than the ground truth due to fewer samples. In contrast, the composite from the DNS with the same computational budget as AI+RES is significantly noisier and contains incorrect physical patterns. For example, the 3-day-lagged Z500 anomaly north of Scandinavia in this composite is a spurious teleconnection likely due to sampling error.
Similarly, Fig. S5 in the SI shows that AI+RES generates physically realistic samples over Chicago that match those from ground truth, while DNS with the same computational budget produces samples with spurious features, obscuring the underlying dynamics.
Discussion
To address the challenge posed by the scarcity of data for studying extreme events, we introduced AI+RES, a novel framework that synergistically combines RES, AI emulators, and GCMs such that they complement each other and alleviate each individual approach’s shortcomings. AI+RES solves the long-standing problem of finding an appropriate score function for RES in a high-dimensional chaotic system. As the first proof-of-concept, we applied this framework on mid-latitude heatwaves in the PlaSim GCM. We showed that this approach enables the sampling of extremely rare events and yields accurate estimates of their return periods up to 50,000 years with an ensemble size of only 400, with a reduction in computational cost of several hundred compared with direct sampling. We also showed that AI+RES significantly outperforms the current state-of-the-art RES that relies on simple score functions.
Our work motivates several avenues for further improvement. First, the deterministic nature of the current emulator limits the quality of its ensemble forecasts. Having an emulator trained directly in a probabilistic manner [10, 45, 46, 47, 48] could enhance the ability of the algorithm to explore multiple plausible pathways to extreme events, improving sampling efficiency and reducing estimator variance. Second, while our current AI emulator focuses exclusively on atmospheric variables, surface temperature extremes are also strongly influenced by land surface processes, particularly soil moisture [49, 50, 51, 44]. Developing a coupled land–atmosphere emulator remains a challenge, but incorporating soil moisture information, potentially through hybrid strategies that combine emulated forecasts with soil moisture state, could yield a more effective score function.
Beyond the specific case of mid-latitude heatwaves, the potential applications of AI+RES are broad. It could easily be applied to other types of weather extremes—such as tropical cyclones, blocking events, precipitation, wind, or compound events. AI+RES offers a powerful tool to generate targeted catalogs of rare events in poorly sampled regions of the state space, particularly where direct sampling of computationally expensive GCMs offers insufficient coverage. The framework could also be well-suited to characterize the tails of distributions in sub-seasonal to seasonal ensemble forecasts, where large ensembles are needed. Moreover, by applying AI+RES to climate models running under different climate change scenarios (e.g., CMIP6 or the emerging global high-resolution runs), it would be possible to explore the evolution of the statistics of rare extremes in future climates. Note that the AI+RES can be readily applied to any state-of-the-art GCM; here, we focused the proof-of-concept on PlaSim as its computational efficiency allows generating a ground truth dataset.
Finally, while this study focused on weather extremes, the approach is fully general. AI+RES could be extended beyond climate science, to any high-dimensional dynamical system for which an AI surrogate model can be constructed.
Code and Data availability
The code implementing AI+RES, along with model configurations, analysis scripts, and the data files used to generate the figures in this paper, will be made publicly available on the project’s GitHub repository upon acceptance. For practical reasons, the full long PlaSim runs will not be publicly posted, but they can be made available upon request from the authors.
Acknowledgments
A.L. was funded in part by RTE France, the French transmission system operator, and benefited from the national agency for research and technology (ANRT) funding under the CIFRE contract 2023/0506. A.W. is supported by the Eric and Wendy Schmidt AI in Science Fellowship, a program of Schmidt Sciences. A.W. and A.L. are also supported by France-Chicago Center FACCTs Award 300004248. This work was also supported by the Institut des Mathématiques pour la Planète Terre (IMPT) and funded in part by the US National Science Foundation through award NSF RISE–2425898 and NSF-2531264. In addition, the authors thank Alessandro Lovo for fruitful discussions and insightful ideas on efficiently coupling RES and AI forecasts, which were invaluable to this project, Anna Asch for her work in making the diagnostic plots of the AI emulator, and Justin Finkel for engaging and insightful discussions. Computational resources were provided by NSF ACCESS (allocation ATM170020), NCAR’s CISL (allocation URIC0009), and the University of Chicago Research Computing Center.
Author contributions
In alphabetical order. D.S.A., F.B., P.H., A.L., C.L.P., J.W. and A.W. conceptualized the work. D.S.A, F.B., P.H., and J.W. supervised and managed the project. A.L. and A.W. curated the data. A.L. and A.W. developed the codes. A.L. conducted the experiments and did the analysis; A.W. led the development of the AI emulator. A.L. and A.W. wrote the original draft. D.S.A., F.B., L.D., P.H., A.L., C.L.P., J.W. and A.W. reviewed and edited the paper.
METHODS
GCM and Data
PlaSim [40] is an efficient general circulation model (GCM) of intermediate complexity that includes a spectral dynamical core that solves the primitive equations for vorticity, divergence, temperature, and humidity. We use PlaSim in this study because it allows rigorous validation of our methodology with large, ground-truth ensemble simulations and it has previously been employed in various studies as a benchmark model to explore the statistics of persistent heat extremes [18, 22], as well as for testing deep learning-based forecasting approaches for such extremes [44, 52]. PlaSim is coupled with simplified representations of land, ocean, and sea ice boundary layers. We use a T42 spectral truncation mapped onto a $64\times 128$ Gaussian grid, with 10 vertical levels. We run PlaSim with sea surface temperature and sea ice thickness fixed to a climatological, annually repeating cycle derived from the AMIP-II boundary dataset (1870–2006) [53], via linear interpolation of monthly climatologies. In the land model, surface temperature is computed through a linear energy balance approach, and soil hydrology is represented using a bucket model with regionally varying water holding capacity[54].
AI Weather Emulator
We train a deep neural network-based AI emulator to predict the full atmospheric state of the PlaSim GCM a short time $\Delta t$ in the future given the current atmospheric state $X_{\text{atm},t}$ and known boundary conditions $X_{\text{bnd},t}$:
$\tilde{X}_{\text{atm},t+\Delta t}=\mathcal{F}_{\Theta}(X_{\text{atm},t},X_{\text{bnd},t}).$ (2)
where $\Theta$ are the trainable parameters of the neural network $\mathcal{F}_{\Theta}$. After initialization, the AI emulator may be cycled autoregressively to obtain long-term predictions. While the AI emulator is trained as a deterministic forecast model, we can generate ensemble forecasts by perturbing the initial conditions.
Architecture
We adopted the PanguWeather 3D Earth-specific transformer (EST) architecture for our AI emulator, which is specifically tailored for global geophysical fields by incorporating Earth-specific inductive biases such as spherical positional encoding and longitudinal periodicity. This architecture has demonstrated strong skill in medium-range forecasting as well as stability in long-term autoregressive simulations [55]. To adapt our emulator to the PlaSim GCM, we made a number of modifications to the original EST used in PanguWeather. To account for the decrease from PanguWeather’s $0.25^{o}$ horizontal resolution to Pangu-PlaSim’s $2.8^{o}$, we decreased the horizontal patch size from $(4,4)$ to $(2,2)$ and the horizontal attention window size from $(6,12)$ to $(2,4)$. We additionally increased the vertical attention window size to include all tokens to account for the inclusion of the top-of-atmosphere boundary variables. The original patch recovery layer was replaced with a pixel shuffle deconvolution layer, as in the ArchesWeather model of [47], to mitigate artifacts arising from patch recovery, while an additional convolutional layer for each variable type (3D prognostic, 2D prognostic, and 2D diagnostic) was added following patch recovery to allow for additional processing of the full-resolution recovered variables.
Training and Validation
Prior to input into the AI emulator, the atmospheric fields from the PlaSim GCM are vertically interpolated from the original 10 topography-following sigma levels to 13 equipressure surfaces. The details of these prognostic atmospheric variables, as well as the other variables input and output by the AI emulator, can be found in Extended Data Table 1. The AI emulator is trained to minimize a weighted sum of mean absolute errors between the model output and the prognostic and diagnostic variables at a timestep $\Delta t=6$ hours in the future. The precise weighted loss function is similar to that used in the PanguWeather model:
$\mathcal{L}(\tilde{X}_{\text{atm},t},X_{\text{atm},t})=\frac{1}{N_{a}N_{z}}\sum_{a,z}\left\|\tilde{X}_{a,z,t}-X_{a,z,t}\right\|_{1}+\frac{1}{4N_{s}}\sum_{s}\left\|\tilde{X}_{s,t}-X_{s,t}\right\|_{1}+\frac{1}{4N_{d}}\sum_{d}\left\|\tilde{X}_{d,t}-X_{d,t}\right\|_{1},$ (3)
where, $X_{a,z,t}$ denotes the value of an atmospheric variable $a$ at pressure level $z$ at time $t$, $X_{s,t}$ and $X_{d,t}$ denote, respectively, the values of surface and diagnostic variables $s$ and $d$ at time $t$, and $\tilde{X}_{\text{atm},t}=\mathcal{F}_{\Theta}(X_{\text{atm},t-\Delta t},X_{\text{bnd},t-\Delta t})$. Each subscripted $N_{\circ}$ denotes the total number of variables of the corresponding type. The net effect of the addition of the normalized MAE for each variable type and the factor of $1/4$ multiplying the surface and diagnostic variable error is that surface and diagnostic variable error is weighted more highly in the loss relative to each atmospheric variable at a particular pressure level.
To train the AI emulator, we use data from the final 100 years of a single 112-year-long PlaSim control run, and reserve year 11 for validation. Prior to input, this data is standardized using statistics from the 100-year training data set so that each variable at each pressure has a global mean of 0 and global standard deviation of 11. We additionally add a Gaussian noise vector independently sampled from $\mathcal{N}(0,2.5\times 10^{-3})$ to each standardized atmospheric variable input to the emulator during training. This noise addition mitigates an instability that can arise during autoregressive prediction with the AI emulator due to the reduced spectral resolution of the PlaSim GCM’s dynamical core relative to its spatial resolution. All other AI emulator training parameters and methods are taken from standard methods for training large-scale vision transformer-based models (see SI Table S1).
Ensemble Forecasting
Ensemble forecasts are generated by adding independent initial condition perturbations to all atmospheric variables at the beginning of the forecast. Perturbations are sampled from a Gaussian distribution with mean 0 and a standard deviation such that the resulting ensemble spread over time is similar to that of the PlaSim GCM with our selected magnitude of perturbations to the initial surface pressure (see SI Section S2).
Rare event Algorithm: Diffusion Monte Carlo
We use a version of the Diffusion Monte Carlo (DMC) algorithm that closely follows the formulations proposed in [19] and [20]. It relies on a one-dimensional score function (or reaction coordinate), $\theta:\mathbb{R}^{d}\to\mathbb{R}$, which assigns higher values to regions of phase space associated with the rare event of interest. The algorithm enhances sampling in regions where $\theta$ is high (via duplication or “splitting”) and suppresses sampling where it is low (via termination or “killing”).
The mathematical properties of DMC have been rigorously studied—see, for instance, [56]—with results establishing its convergence and asymptotic behavior as the ensemble size NN tends to infinity. Under mild integrability assumptions, DMC provides unbiased estimators that converge in the large-NN limit. These convergence results are quite general and remain valid even for systems of arbitrarily high dimension $d$. Any statistical quantity that can be computed via direct sampling can, in principle, also be estimated using DMC. This includes observables that depend on the full trajectory of the process from the initial time $t_{0}$ up to a given time $t_{f}$. We denote by $\left(X_{t}^{n}\right)_{1\leq n\leq N}$ the NN realizations generated during the simulation and refer to each sample $X_{t}^{n}$ as a walker.
We introduce a finite sequence of scheduled resampling times $0=t_{0}<t_{1}<\cdots<t_{K}=t_{f}$, which may be spaced either uniformly or non-uniformly. For notational convenience, we set $X_{k}^{i}:=X_{t_{k}}^{i}$. The resampling steps are governed by a sequence of splitting functions $V_{k}$, which are themselves functions of the score function $\theta$.
The procedure is described in Algorithm 1.
Algorithm 1 Diffusion Monte Carlo algorithm
Input: NN walkers $\left(X_{0}^{n}\right)_{1\leq n\leq N}$ starting from an initial condition.
Initialize: Choose a sequence of resampling times $0=t_{0}<t_{1}<\cdots<t_{K}$ and a family of splitting functions $(V_{k})_{k}$ depending on a score function $\theta$.
1:for $k=0$ to $K$ do
2:  (i) Reweighting: For each walker $i$
3:  if $k=0$ then
4:   Define initial weights: $w_{0}^{i}=\exp\left(V_{0}(X_{0}^{i})\right)$
5:  else
6:   Define weights: $w_{k}^{i}=\bar{w}_{k-1}\exp\left(V_{k}(X_{k}^{i})-V_{k-1}(\hat{X}_{k-1}^{i})\right)$
7:  end if
8:  Compute average weight: $\bar{w}_{k}=\frac{1}{N}\sum_{i=1}^{N}w_{k}^{i}$
9:  (ii) Resampling: Create an updated ensemble of walkers $\left(\hat{X}_{k}^{i}\right)_{1\leq i\leq N}$ by sampling $N_{k}^{i}$ copies of each $X_{k}^{i}$ such that $\sum_{i}N_{k}^{i}=N$ and $\mathbb{E}[N_{k}^{i}]=\frac{w_{k}^{i}}{\bar{w}_{k}}$.
10:  (iii) Simulation: Integrate the model from $t_{k}$ to $t_{k+1}$: $\hat{X}_{k}^{i}$ $\rightarrow$ $X_{k+1}^{i}$.
11:end for
Note: walkers for which $N_{k}^{i}=0$ are terminated and replaced by copies of walkers with $N_{k}^{j}\geq 2$.
One should note that the DMC algorithm applies a selection procedure before the first simulation step. In practice, if one does not want to apply this first selection step, it can be avoided by choosing $V_{0}=0$. We do this in this study.
We perform the resampling step $(ii)$ using the pivotal sampling scheme [57]. For a stochastic model, duplicated walkers will separate naturally after resampling. For the deterministic chaotic system we work with, it is necessary to perturb the model immediately after each resampling step. Following the approach of [18], we apply perturbations to the spherical harmonics of the logarithm of the surface pressure field. In our implementation, we use a perturbation amplitude of $3\times 10^{-3}$.
For any function $\psi$ of the state (or history) of the system, the DMC algorithm yields the following unbiased estimator:
$\mathbb{E}\left[\psi\left(X_{k}\right)\right]\approx\frac{\bar{w}_{k-1}}{N}\sum_{i=1}^{N}\psi\left(X_{k}^{i}\right)e^{-V_{k-1}\left(\hat{X}_{k-1}^{i}\right)}$ (4)
where the expectation on the left is with respect to the distribution of the original process without RES. Using this formula with the choice $\psi=\mathbbm{1}_{A\geq a}$ for a large $a\in\mathbb{R}$ and an observable AA of the trajectory of $X$, one then has access to the probability of AA taking extreme values $\left(\mathbb{E}\left[\mathbbm{1}_{A\geq a}\right]=\mathbb{P}\left[A\geq a\right]\right)$. In our case, this observable is $A_{L}(t_{f})$ defined as the $L$-day average of the 2m temperature over the region $\mathcal{R}$ in Eq. (1).
Following [20], we choose a splitting function of the form $V_{k}=C_{k}\theta(X_{k})$, where $C_{k}>0$ is a constant that can vary at each resampling step. The choice of $C_{k}$ is critical as it controls the strength of the selection process; we discuss it in the Supplementary Information, Section S4.3.
To improve the algorithm’s robustness to the choice of $V_{k}$, [19] introduced an additional step known as quantile mapping, which gives the algorithm its full name: Quantile Diffusion Monte Carlo (QDMC). Further details can be found in the Supplementary Information, Section S4. In our work, however, we adopted a simpler alternative by rescaling the score function $\theta$ at each resampling step using the following transformation
$\tilde{\theta}_{k}(X^{i}_{k})=\frac{\theta(X^{i}_{k})-\mu_{\theta_{k}}}{\sigma_{\theta_{k}}},$ (5)
where $\mu_{\theta_{k}}$ and $\sigma_{\theta_{k}}$ denote the mean and standard deviation of $\theta$ across all walkers at time $t_{k}$. The rescaled score function $\tilde{\theta}_{k}$ is then used to compute the splitting function via $V_{k}=C_{k}\tilde{\theta}_{k}$. Using this approach, we obtained results that were very similar to those achieved with the quantile-mapping step.
Choice of the score function
At each resampling time and for each walker $i$, we perform an ensemble forecast from the AI emulator with $M$ members starting from the last state of the walker, until the target period $t_{f},t_{f}+L]$. This yields an ensemble of M forecasts ${\hat{A}^{i,j}_{L}(t_{f}|X^{i}_{k})}_{j=1}^{M}$ of the observable $A^{i}_{L}(t_{f})$ (see Eq. (1)) starting from the initial condition $X^{i}_{k}$.
We combine these ensemble forecasts into the score function by taking the ensemble mean of the forecast observable $\hat{A}^{i,j}_{L}(t_{f}|X^{i}_{k})$:
$\theta(X^{i}_{k})=\frac{1}{M}\sum_{j=1}^{M}\hat{A}^{i,j}_{L}(t_{f}|X^{i}_{k})$ (6)
We set the number of ensemble members per forecast to $M=100$. Since our score function is defined using the ensemble mean, we do not expect substantial performance gains from increasing $M$. This parameter could, however, play a more significant role under alternative formulations of the score function.
We note that $A^{i}_{L}(t_{f})$ could be predicted by other methods. We choose to leverage the AI emulator because of its proven forecasting skill and numerical efficiency, but other methods could be used. In particular, it would be interesting to test simpler methods based on directly learning the observable $A_{L}(t_{f})$ (or its distribution) from an initial condition, such as in [58, 59, 44, 60, 61, 62]. In this work, we also test the best possible forecast system, using ensembles of the PlaSim model itself to generate the score function as described in the Different baselines section.
We note also that one could use a different statistic calculated from the ensemble forecast as the score function. The ideal score function to estimate the probability of achieving the rare event $A_{L}(t_{f})>a$ for large $a$ should anticipate walker paths leading to the rare event. An alternative score function appropriate for this goal is the conditional probability function $\theta(x)=P(A_{L}(t_{f})>a|X_{t_{k}}=x)$, which is referred to as the committor function [63]. This score function has been proven to be optimal for a splitting algorithm similar to quantile DMC [64]. However, in our case there is no clear threshold, $a$, since we aim to characterize the entire tail of the distribution of $A_{L}(t_{f})$. Our choice of score function, namely, the mean of the ensemble forecasts, favors the duplication of walkers for which the conditional expectation $\mathbb{E}[A_{L}(t_{f})\mid X_{t_{k}}=x]$ is large. This approach is more conservative than selecting a fixed threshold a priori.
Choice of other parameters
The important parameters of the algorithm are given in Extended Data Table 2. We use $K=6$ steps with scaling constants $(C_{1},C_{2},C_{3},C_{4},C_{5},C_{6})\approx(0.,0.,0.,1.6,1.8,2.0)$, tuned empirically. The steps with $C_{k}=0$ ensure that the walkers are well-separated in phase space when selection begins.
Return period curves
To assess the ability of our model to accurately estimate the probability of very rare events, a classical diagnostic is the return-period curve. Since we have constrained the target observable to a specific period in summer, we only observe a single event per year (i.e., per summer simulated). In this setting, the return period $T_{a}$ associated with a return value $a$ is defined as the inverse of the probability of exceedance $p_{a}=\mathbb{P}(A_{L}(t_{f})>a)$. A return-period curve is then defined as the plot of the return values $a$ as a function of their corresponding return periods $T_{a}$.
Let $(a_{1},a_{2},\ldots,a_{N})$ be all the values of the observable $A_{L}(t_{f})$ sampled by DNS or RES. To draw the return-period curve, we first need to estimate $p_{a_{j}}$ for each value $a_{j}$. For DNS, we estimate the probability of exceedance with the following empirical estimator:
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(Y_{t_{f}+T}^{i}\right)$ (7)
where, in this expression, $Y_{t}^{i}$ are independent PlaSim simulations without RES.
For the RES algorithm, we use the formula (4) with the observable $\psi(X)=\mathbbm{1}_{A_{L}(t_{f})>a_{j}}$ to compute the probability of exceedance. Namely,
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{\bar{w}_{K}}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(X_{t_{f}+T}^{i}\right)e^{-V_{K}\left(\hat{X}_{K}^{i}\right)}$ (8)
The return period for each value is then obtained by taking the inverse of its probability of exceedance.
Extreme value theory fits
To compare the uncertainty of our return period estimates from RES with those obtained via Extreme Value Theory (EVT), we use the standard Peak-Over-Threshold (POT) method, fitting a Generalized Pareto Distribution (GPD) to the data.
The family of distributions $GPD(\sigma,\xi)$ is suitable to characterize the tail of random variables. If $V_{1},\dots,V_{n}\in\mathbb{R}$ is a sequence of independent and identically distributed random variables, then, under suitable assumptions, the asymptotic distribution of the excesses Z=Vu|V>uZ=V-u\ |\ V>u over a sufficiently high threshold $u$ is given by:
$F_{(\sigma,\xi)}(z)=1-\left[1+\xi\frac{(z-u)}{\sigma}\right]^{-1/\xi}_{+},$ (9)
where $a]_{+}=\max\{0,a\}$, $\sigma>0$ is the scale parameter, $\infty<\xi<\infty$ is the shape parameter and $F_{(\sigma,\xi)}$ is the cumulative distribution function. The parameters $(\sigma,\xi)$ are then fitted on the exceedances $Z$ and used to determine the GPD return values.
To ensure a fair comparison in terms of sample size, the GPD distributions were fitted on datasets of the same size as the AI+RES experiments (namely $N=400$ in Fig. 2), on 100 independent training datasets. The gray shading represents the 10th–90th percentile range of return-period curves obtained from the GPD fits across these datasets.
In the Peak-Over-Threshold method, the key hyperparameter to tune is the threshold $u$ above which the data is modeled. We determine $u$ using the heuristic of the Mean Residual Life plot ([65]), selecting a percentile that can be applied consistently across all datasets. Based on this criterion, we retain the 90th percentile of the data as the threshold. Parameters were estimated using the scipy Python package ([66]).
Derivation of the computational speed-up factors
In the following, we derive two criteria to evaluate the computational gains provided by the AI+RES algorithm compared to DNS. We make the key assumption that the computational cost of running the AI emulator ensemble forecasts is negligible compared to that of running the physics-based model (in this case, the GCM). For reference, [7] report speed-ups of up to $\mathcal{O}(10^{4}\!-\!10^{5})$ and energy-consumption reductions of $\mathcal{O}(10^{4})$ for FourCastNet (AI emulator) relative to the IFS model (numerical weather prediction model). In our case, the cost comparison between the AI emulator and PlaSim is less favorable, as PlaSim is specifically designed to be computationally inexpensive. This is precisely what allows us to run a 50,000-member PlaSim ensemble within reasonable computational limits, enabling a robust validation of the methodology. Nevertheless, a preliminary analysis indicates that our emulator runs approximately 10 times faster on a single A100 GPU than PlaSim using 64 Intel Xeon Platinum 8160 CPU cores, even without any model variable or weight pruning, or inference-specific optimizations. This does not undermine our assumption of negligible cost for the AI emulator, as the present study serves as a proof of concept.
Sampling speed-up factor
One of the goals of RES is to efficiently generate a large catalog of extreme events. To quantify the improvement achieved by RES in this task, we count the average number of events, $n_{a}$, that the algorithm samples above a given threshold $a$ (corresponding to a return period $T_{a}$), using a fixed budget of $N_{RES}$ walkers.
To obtain the same number of events with DNS, one would need to have on average a budget of $N_{DNS}(a)=T_{a}n_{a}$ members by definition of the return period. We then introduce the Sampling Speed-up Factor (SSUF) as the ratio between the DNS and RES budgets needed to sample the same number of extreme events:
$\text{SSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{T_{a}n_{a}}{N_{RES}}$ (10)
We also introduced a metric to quantify the diversity of the sampled trajectories (Section S7 of the SI), demonstrating that the generated samples are not overly correlated.
Variance speed-up factor
The other main objective of RES is to reduce the uncertainty in rare probability estimates. Following [19, 20], we also quantify speed-up based on the reduction in the variance of estimated rare event probabilities. Suppose we have used RES to obtain an estimate of the probability of exceedance $\hat{p}_{a}$ above the value $a$ with a variance $\sigma_{RES}(a)^{2}$ using a budget of $N_{RES}$ walkers. We can compare this with the variance of the DNS estimator (Eq. (7)) as a function of NN and $a$, which is given by
$\sigma_{DNS}(a,N)^{2}=\frac{p_{a}(1-p_{a})}{N}$ (11)
We can then estimate the number of DNS samples that would be necessary to produce a similarly accurate estimate as RES as follows
$N_{DNS}(a)\approx\frac{\hat{p}_{a}(1-\hat{p}_{a})}{\hat{\sigma}_{RES}(a)^{2}}.$ (12)
We then define the Variance Speed-up Factor (VSUF) as
$\text{VSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{p_{a}(1-p_{a})}{\hat{\sigma}_{RES}(a)^{2}N_{RES}}=\frac{T_{a}-1}{T_{a}^{2}\hat{\sigma}_{RES}(a)^{2}N_{RES}}$ (13)
In practice, we estimate the variance of RES, $\hat{\sigma}_{RES}(a)^{2}$, empirically from only 10 independent realizations of the algorithm, as running a larger number of experiments is computationally expensive. Consequently, the estimate of $\hat{\sigma}_{RES}(a)^{2}$ itself is subject to sampling error. In Section S8 of the SI, we also present alternative estimates of the variance using only a single realization of the algorithm.
[…]
Supplemental Information for:
AI-boosted rare event sampling to characterize extreme weather
Amaury Lancelin, Alex Wikner, Laurent Dubus, Clément Le Priol, Dorian S. Abbot, Freddy Bouchet, Pedram Hassanzadeh, and Jonathan Weare
S1 Why traditional rare event sampling methods fail
A large part of the applications of trajectory splitting-based RES algorithms in climate science has focused on sampling large deviations of long-time averages of an observable AA [18, 21, 22, 23], where the averaging window typically spans several months. In these approaches, the score function is usually constructed from instantaneous values of AA (or from its average over the most recent resampling interval) at the resampling time. While this strategy is effective for sampling long-time averages, it is not well-suited for sampling extreme events which occur on time scales shorter than the Lyapunov time of the system—about 5–10 days in the atmosphere. To address such events, a score function capable of anticipating extremes weeks in advance is required. This is precisely the approach taken in the present work, where forecasts from an AI-based emulator of the climate model are leveraged to construct the score function.
S2 How we generate PlaSim ensembles
As described in the Methods section, for the duplication of walkers at each resampling step in the DMC algorithm to have an effect, it is necessary to perturb their initial conditions. Following [18], we introduce perturbations by adding a small noise to the coefficients of the spherical harmonics of the logarithm of the surface pressure field. Fig. S1 shows the time evolution of the standard deviation of 100-member PlaSim ensembles generated with different values of $\epsilon$. In our DMC implementation, we use a perturbation amplitude of $\epsilon=3\times 10^{-3}$. We tested several values of $\epsilon$ and found that this choice provides a good balance: it enhances the ability of the PlaSim ensembles to explore the phase space, while remaining small enough to avoid altering the final statistics of the target observable or introducing non-physical discontinuities in the trajectories. We anticipate that more sophisticated perturbation strategies, such as Bred vectors [67] or early splitting [30], could further improve the performance of the algorithm.
FIG. S1: Time evolution of the standard deviation of PlaSim ensembles for different values of perturbation amplitude $\epsilon$. The perturbation is applied to the spherical harmonic coefficients of the logarithm of the surface pressure field of the initial condition. The ensemble size is 100 and the results are averaged over 10 independent initial conditions. Top panels: Standard deviation of the spatially-averaged daily T2m for the France (left) and Chicago (right) regions. Bottom panels: Standard deviation of the spatially-averaged daily $Z_{500}$ for the France (left) and Chicago (right) regions. The black, blue, and red curves show results for $\epsilon=10^{-4}$, $\epsilon=10^{-3}$, and $\epsilon=10^{-2}$, respectively.
S3 The Pangu-PlaSim Emulator
S3.1 Training hyperparameters
Table S1: Hyperparameters used to train the Pangu-PlaSim emulator.
Hyperparameter Value
Optimizer AdamW
Learning rate scheduler OneCycleLR
Total epochs 100
Annealing epochs 10
Min. learning rate 1e-6
Max. learning rate 1e-4
Batch size 64
Weight decay 3e-6
Drop path rate 0.2
Epoch selection Lowest loss
S4 Choosing the hyperparameters of the algorithm
S4.1 Quantile Diffusion Monte Carlo (QDMC)
In the Diffusion Monte Carlo (DMC) method, the splitting functions $\left(V_{k}\right)_{k}$ can critically influence the behavior of the algorithm. A frequently used choice is $V_{k}(X)=C_{k}\theta(X)$, where $C_{k}>0$ controls the duplication rate of walkers. While simple, this formulation has drawbacks. In nonlinear systems, walkers with large $\theta$ values may produce an excessive number of offspring, eventually leading to the so-called extinction phenomenon in which the population becomes overly concentrated around a few trajectories. Although such a scheme remains unbiased, the variance of the resulting estimator can be prohibitively large. On the opposite end, if selection is too weak, particle weights are nearly equal, and the method essentially reduces to naive Monte Carlo sampling—again yielding high variance. Proper tuning is therefore essential to balance these two extremes.
To alleviate these issues, [19] proposed a modification of DMC designed to improve robustness. The idea is to rescale the score function $\theta$ dynamically at each resampling step so that its distribution matches a prescribed target distribution $\nu_{k}$, most often taken as Gaussian. The transformed score, denoted $\theta_{k}^{\prime}$, is then used in the splitting and pruning procedure. This approach is known as Quantile Diffusion Monte Carlo (QDMC).
The distinguishing feature of QDMC is this rescaling step. After estimating the empirical distribution of $\theta(X_{t_{k}})$, one constructs a transformation $\theta_{k}^{\prime}=\gamma_{k}(\theta)$ such that the distribution of $\theta_{k}^{\prime}(X_{t_{k}})$ is close to $\nu_{k}$. More precisely, QDMC introduces a transport map from the empirical law of $\theta_{k}$ to the target distribution $\nu_{k}$ via
$\gamma_{k}(y)=F_{\nu_{k}}^{-1}(F_{\theta_{k}}(y)),$ (S1)
where $F_{\theta_{k}}$ is the cumulative distribution function (CDF) of $\theta_{k}$, and $F_{\nu_{k}}^{-1}$ is the quantile function (inverse CDF) of $\nu_{k}$, typically $\mathcal{N}(0,1)$.
An attractive property for rare-event simulation algorithms is invariance with respect to monotone bijective transformations of the score function $\theta$. Unlike standard DMC, QDMC satisfies this invariance, which provides greater flexibility. Nevertheless, some elements of the method, such as the choice of target distribution or the splitting constants $C_{k}$, remain somewhat arbitrary and require further discussion.
In our experiments, we observed that when the number of walkers NN is too small, the quantile-mapping step—based on the empirical estimate of $F_{\theta_{k}}$ which only relies on the ranking between walkers—can have an undesired effect, sometimes assigning very different weights to walkers with very close scores. To avoid this issue, we opted for a simpler normalization procedure described in Eq. (5) in the main text.
S4.2 Choice of the resampling scheme
Algorithm S1 Pivotal Resampling (adapted from [57])
Input: Normalized weights $\{\tilde{w}_{k}^{1},\dots,\tilde{w}_{k}^{N}\}$ with $\sum_{i=1}^{N}\tilde{w}_{k}^{i}=N$.
Output: Number of clones $N_{k}^{i}$ for each particle $i=1,\dots,N$ at resampling step $k$.
1:Step 1: Decomposition. For each $i$, decompose
$\tilde{w}_{k}^{i}=\lfloor\tilde{w}_{k}^{i}\rfloor+\delta_{k}^{i},\qquad\delta_{k}^{i}\in[0,1).$
2:Step 2: Pivotal sampling on fractional parts.
3:while there exist at least two indices $i,j$ with $\delta_{k}^{i},\delta_{k}^{j}\in(0,1)$ do
4:  Let $s=\delta_{k}^{i}+\delta_{k}^{j}$.
5:  if $s\leq 1$ then
6:   With probability $\frac{\delta_{k}^{i}}{s}$: set $\delta_{k}^{i}\leftarrow s,\;\delta_{k}^{j}\leftarrow 0$.
7:   Otherwise: set $\delta_{k}^{j}\leftarrow s,\;\delta_{k}^{i}\leftarrow 0$.
8:  else
9:   With probability $\frac{1-\delta_{k}^{j}}{2-s}$: set $\delta_{k}^{i}\leftarrow 1,\;\delta_{k}^{j}\leftarrow s-1$.
10:   Otherwise: set $\delta_{k}^{j}\leftarrow 1,\;\delta_{k}^{i}\leftarrow s-1$.
11:  end if
12:end while
13:Step 3: Final number of clones.
14:for $i=1$ to NN do
15:  $N_{k}^{i}\leftarrow\lfloor\tilde{w}_{k}^{i}\rfloor+\delta_{k}^{i}$
16:end for
17:return $\{N_{k}^{i}\}_{i=1}^{N}$
As described in Algorithm 1 in the main text, at each resampling step $k$, we duplicate each walker a random number of times $N_{k}^{i}$, subject to the constraints
$\mathbb{E}[N_{k}^{i}]=\frac{w_{k}^{i}}{\bar{w}_{k}}\coloneq\tilde{w}_{k}^{i},\quad\sum_{i=1}^{N}N_{k}^{i}=N.$ (S2)
Several resampling strategies can be used to generate the integers $N_{k}^{i}$ while preserving these properties. In this work, we employ the pivotal resampling scheme [57], detailed in Algorithm S1, which we found yields lower variance than alternative approaches.
S4.3 Scaling the constants $C_{k}$
For the splitting functions, we chose to set $V_{k}=C_{k}\theta_{k}$, where $C_{k}>0$ is a constant to tune. A higher value of $C_{k}$ results in a larger number of walker clones, making the algorithm more selective. However, if $C_{k}$ is too large, it may lead to the extinction phenomenon described in Section S4.1, whereas if it is too small, the algorithm behaves similarly to direct numerical sampling by not sampling enough extreme events. It is often desirable not to use the same constant $C_{k}$ for all resampling steps. In our case, we are forecasting the future state of the trajectory to compute the score function, and the uncertainty in the forecast decreases as we approach the final time $t_{f}$. Therefore, we chose to set $C_{k}=Ce^{-\alpha(t_{f}-t_{k})}$, where $C>0$ is a constant and $\alpha>0$ is a decay factor. A similar approach is used in [19]. Both $C$ and $\alpha$ are hyperparameters of the algorithm. Using a non-exhaustive grid search, we found that setting $C=2.0$ and $\alpha=0.02$ yielded good results in terms of variance of the final estimators of rare event probabilities. Furthermore, we turned off resampling at the first three resampling times, $t_{1}$, $t_{2}$, and $t_{3}$, by setting $C_{1}=C_{2}=C_{3}=0$ to ensure that the walkers were sufficiently separated in phase space before actually applying a resampling.
While we expect that further tuning of the hyperparameters could improve performance, determining the optimal values for all hyperparameters involved in the algorithm is computationally expensive and, although mathematically interesting, lies beyond the scope of this work.
S4.4 Choice of the initial condition
Motivated by sampling the most extreme events possible, we chose for the France region an initial condition at $t_{0}$ corresponding to the largest values of the observable $A_{L}(t_{f})$ (see Eq. (1) in the main text) in the 100-year training set of the emulator. For the Chicago region, in order to test the robustness of the algorithm to the choice of initial condition, we instead selected a random initial condition at $t_{0}$ from the same training set.
S5 Computing error bars for the DMC algorithm
To compute variance estimates with the DMC algorithm, we performed 10 independent runs of the algorithm with the same hyperparameters and the same initial conditions. We then computed the empirical variance of the return period estimates across the 10 runs. The same methodology is used in [22, 18, 29].
However, it is possible to compute variance estimates for the DMC algorithm using a single realization of the algorithm [68, 19, 20]. As discussed in [20], two different approaches can be used to estimate the variance in DMC: one that systematically overestimates it, and another that underestimates it. For a general DMC estimator $\hat{f_{\psi}}$ of the quantity $\mathbb{E}[\psi(X_{k})]$, defined via Eq. (4) in the main text, the pessimistic variance estimate, which leans toward overestimation, is given by
$\hat{\sigma}^{2}_{\mathrm{pess}}=\frac{1}{N^{2}}\sum_{i=1}^{N}\left|\sum_{\mathrm{anc}(X_{k}^{j})=i}\psi\left(X_{k}^{j}\right)\bar{w}_{k-1}e^{-V_{k-1}\left(\hat{X}_{k-1}^{j}\right)}\right|^{2}-\hat{f_{\psi}}^{2}.$ (S3)
Here, $\mathrm{anc}(X_{k}^{j})$ identifies the index of the original ancestor at $t=t_{0}$ from which the trajectory $X_{k}^{j}$ descends. Conversely, an optimistic estimator, which tends to underestimate the variance, is expressed as
$\hat{\sigma}^{2}_{\mathrm{opt}}=\frac{1}{N^{2}}\sum_{i=1}^{N}\left|\psi\left(X_{k}^{i}\right)\bar{w}_{k-1}e^{-V_{k-1}\left(\hat{X}_{k-1}^{i}\right)}\right|^{2}-\hat{f_{\psi}}^{2}.$ (S4)
The intuition behind these two forms is the following: the optimistic estimator considers each trajectory as an independent contribution, similar to importance sampling [69], while the pessimistic estimator aggregates all descendants of a given ancestor into a single effective data point. In practice, the effective number of independent samples lies between these two limiting cases. As there is no theoretical guarantee that one of these estimators is better than the other, we preferred to use the empirical variance computed from the 10 independent runs of the algorithm to compute variance speed-up factors (see Eq. (13) and Fig. 3 in the main text). In Fig. S2, we verify that the empirical variance lies between the pessimistic and optimistic estimators, within statistical error.
Refer to caption Refer to caption
FIG. S2: Comparison of speed-up factors obtained via the empirical, the pessimistic (Eq. (S3)) and optimistic (Eq. (S4)) variance estimates. The experiments shown here are described in Extended Data Table II in the main text, for Chicago (left panel) and France (right panel). The variance speed-up factor measures the reduction in the variance of return time estimates compared to direct numerical sampling (Eq. (13) in the main text). The blue curve shows the empirical speed-up ratio obtained by running the AI+RES algorithm 10 times and computing the variance of the return time estimates across the 10 runs. The green curve shows the mean over the 10 optimistic variance estimates, each computed from a single realization of the algorithm. The orange curve shows the mean over the 10 pessimistic variance estimates, each computed from a single realization of the algorithm.
S6 Interpretation of the return times
Since we have constrained the target period to a specific moment in summer, we only observe a single event per year (i.e., per summer simulated). In this setting, the return time associated with a return value $a$ is defined as the inverse of the probability of the event $A_{L}(t_{f})>a$.
In our framework, these return times can be interpreted as the average number of years one would have to wait before observing an event of magnitude greater than $a$, assuming we repeatedly simulate the same summer, each time conditioned on the same initial state at the beginning of the season. In that sense, they correspond to conditional return times. Given our finite computational budget, we chose this approach to allow for a larger number of experiments, thereby enabling a robust validation of the algorithm. However, extending this to the unconditional case is straightforward. This can be achieved either by initiating the simulations earlier in spring—so that the influence of initial conditions no longer persists into the target period $t_{f},t_{f}+L]$—or by relying on appropriate heuristics, such as those proposed in [35].
S7 Characterizing the diversity of trajectories sampled with the Rare event algorithm
FIG. S3: Most Recent Common Ancestor Distance (MRCAD) for AI+RES sampled trajectories. The figure shows results for France and Chicago with the setup given in Extended Data Table II in the main text. The MRCAD is the average number of resampling steps since the current population last shared a common ancestor (max = 7, min = 1, with $K=6$ resampling steps). The solid lines show the mean MRCAD over 10 independent realizations of the AI+RES algorithm, with the shading indicating the $95\%$ confidence interval. A larger MRCAD indicates greater trajectory diversity. The AI+RES generated trajectories maintain substantial diversity, yielding a representative catalog of rare events.
The DMC algorithm duplicates walkers at each resampling step to generate a large catalog of extreme events. When two walkers share the same parent, their trajectories are highly correlated, which reduces the diversity of the sampled events. To quantify this diversity, we introduce the Most Recent Common Ancestor Distance (MRCAD). The MRCAD is defined as the average number of resampling steps since a given population $\mathcal{P}$ last shared a common ancestor, with a maximum of 7 and a minimum of 1 for $K=6$ resampling steps. A higher MRCAD indicates greater diversity in the sampled trajectories.
To assess how diversity is affected when focusing on rarer events, we computed the MRCAD for different populations $\{\mathcal{P}_{p}\}_{p}$, where $\mathcal{P}_{p}$ consists of walkers whose associated values of $A_{L}(t_{f})$ have a return time larger than $1/p$. Figure S3 shows the MRCAD as a function of the return period $1/p$ for AI+RES experiments. We observe that trajectory diversity remains relatively high (between 4 and 5) up to return periods of approximately $10^{3}$ years for both regions studied. Beyond this threshold, diversity decreases rapidly for Chicago, whereas it remains high for France. These results indicate that the AI+RES algorithm successfully samples a large catalog of extreme events with sufficient diversity, rather than merely replicating a single trajectory with an extreme value of $A_{L}(t_{f})$.
S8 Using PLASIM itself to produce forecasts
FIG. S4: Return-time curve comparison between AI+RES and PFS+RES. The setup matches that described in Extended Data Table II in the main text, except that it uses $N=200$ walkers (instead of $N=400$), since larger values are computationally prohibitive for PFS+RES. Black dots are the empirical return periods obtained with a $50,000$-member control simulation. The red dots are from this control simulation but using the same computational budget as the AI+RES algorithm ($N=200$). The solid blue line shows the median return time curve produced by 10 independent realizations of the AI+RES algorithm with $N=200$ walkers and the blue-shaded area represents the 10th to 90th percentile range. The solid green line shows the median return time curve produced by 10 independent realizations of the PFS+RES algorithm with $N=200$ walkers and the green shaded area represents the 10th to 90th percentile range. The gray shaded area is obtained by fitting different GDP distributions with independent $N=200$ training datasets, and showing the 10th to 90th percentile range.
The Perfect-Forecast-System+RES (PFS+RES) approach is designed to provide an upper bound on algorithmic performance. As noted earlier, using the AI-emulator introduces an imperfect approximation of the score function due to the emulator’s limited weather forecast skill, particularly beyond 10–15 days of lead time. Here, we propose performing the ensemble forecast directly with the PlaSim GCM, thereby obtaining a near-perfect forecast to use in the score function.
We conducted experiments using the same setup as the AI+RES experiments, with parameters listed in Extended Data Table II in the main text. The only difference is that the number of walkers was set to $N=200$ instead of $N=400$, as the latter is computationally prohibitive for PFS+RES.
Extended Data Fig. 7 of the main text shows the return-time curves for AI+RES and PFS+RES, demonstrating that PFS+RES can produce highly accurate return-period estimates with smaller variance than AI+RES. To quantify precisely how much we lose in terms of variance when moving from a perfect forecast system (PFS+RES) to an imperfect one (AI+RES), we compared the speed-up factors relative to DNS for both methods in Extended Data Fig. 7. Importantly, these speed-up factors assume the cost of running ensemble forecasts is negligible. While this is not the case for PFS+RES, the comparison provides an indication of the potential performance achievable with near-perfect forecasts, under the assumption that forecast computations are inexpensive relative to the cost of running the physical model (here, PlaSim).
S9 Additional results
FIG. S5: Composite maps of heatwaves over Chicago with return periods exceeding 100 years. Events are defined by Eq. (1) of the main text with $L=$ 3 days. The first row shows daily mean Z500 anomaly composites three days before the heatwave onset; the second row shows the $L$-day average Z500 anomaly composites during the heatwave; the third row shows the $L$-day average T2m anomaly composites during the heatwave. First column: DNS with N=50, [^0] N=50{,}000 (ground truth). Second column: results from the AI+RES algorithm with $N=400$. Third column: DNS with $N=400$.
FIG. S6: Correlation between predicted and actual $A_{L}(t_{f})$ across forecast systems and regions. The blue curves correspond to forecasts made with the AI emulator within AI+RES for Chicago, while gold shows AI emulator results for France. Red curves represent forecasts made with the PlaSim GCM itself—referred to as the Perfect Forecast System (PFS) for the France region. Results are based on 10 independent realizations of the experiments in Extended Data Table II in the main text. For PFS, the number of walkers is $N=200$ instead of $N=400$ for the AI+RES experiments, and the analysis is limited to France, for computational reasons. At each resampling step in AI+RES, score functions $\theta_{k}^{i}$ are computed from ensemble forecasts (with $M=100$ members) with the AI emulator for each PlaSim walker (Eq. (6) in the main text). Here, three resampling steps are used (at $t_{f}-t_{k}=10$, 55, and 0 days before the event). We show the correlation between predicted (ensemble mean from the emulator) and actual (PlaSim realizations) $A_{L}(t_{f})$ for all events (left) and for events with return periods >100>100 years (right). For rare events, the emulator forecasts are notably less skillful at the earliest resampling time in France compared to Chicago, likely reflecting the stronger role of soil moisture in the former region (since soil moisture is absent from emulator inputs but evolved by PlaSim). This result explains the higher variance of rare-probability estimators in France compared to Chicago (left panel of Fig. 3 in the main text). Similarly, the smaller variances obtained for PFS-RES compared to AI-RES (Extended Data Fig. 7. of the main text), can also be attributed to an overall higher forecast skill of PFS both for typical and extreme events.